OpenStudy (anonymous):

Given f(x)=x^2+1, determine f(1/x), x does not equal to 0. Why is it:

OpenStudy (anonymous):

well plug in 1/x to the function so thats (1/x)^2 + 1 that equals (1/x^2) + 1 x cannot equal zero because 1/0 is undefined

OpenStudy (anonymous):

Why is the process: \[\left(\begin{matrix}1 \\ x ^{2}\end{matrix}\right) + \left(\begin{matrix}1\times x ^{2} \\ 1 \times x ^{2}\end{matrix}\right)\]

OpenStudy (anonymous):

\[\left(\begin{matrix}1 \\ x ^{2}\end{matrix}\right) + \left(\begin{matrix}x ^{2} \\ x ^{2}\end{matrix}\right)\]

OpenStudy (anonymous):

answer: \[\left(\begin{matrix}1 + x ^{2} \\ x ^{?}\end{matrix}\right)\] Can someone explain to me why it's this process I should do?

OpenStudy (anonymous):

x≠ 0 is just a condition to guarantee f(1/x) exist! You don't have to do anything.

OpenStudy (anonymous):

that's what my teacher showed me but I'm having trouble understanding it.

OpenStudy (anonymous):

x= 0 is not belong to domain of f(1/x) Since if x= 0 --> f(1/x ) = ∞ ( undefined)

OpenStudy (anonymous):

ohh, okay.

OpenStudy (anonymous):

That's the rule, just memorize it whenever you see fraction, domain is numerator must be≠ 0

OpenStudy (anonymous):

alright, thank you!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I mean *denominator* must be ≠ 0!

OpenStudy (anonymous):

okay!